OpenStudy (anonymous):

Passing through (3,-6) and perpendicular to the line whose equation is x+4y=9. What is the slope intercept form? and point slope of the line?

OpenStudy (jit4won):

x/4.5 +y/-18 = 1

OpenStudy (anonymous):

for point slope i get y+6=4x-12 and for slope form i get y=4x-18 is this correct?

OpenStudy (jit4won):

y= 4x-18 is correct

OpenStudy (anonymous):

thanks

OpenStudy (jit4won):

slope intercept form is x/4.5 +y/-18 = 1

OpenStudy (anonymous):

okay